OpenStudy (anonymous):

In the triangle below, what ratio is csc G? (i really just need the rule for csc, so i can figure out the other problems on my own. they are pretty easy)

hero (hero):

\[\csc(\theta) = \frac{1}{\sin(\theta)} = \frac{\text{hypotenuse}}{\text{opposite}}\]